OpenStudy (anonymous):

find the product of the complex number and its conjugate (-1)-(squareroot 2i)

hartnn (hartnn):

\((a+bi)(a-bi)=a^2+b^2\) does this help?

OpenStudy (anonymous):

kinda how do you use it?

hartnn (hartnn):

u have \(-1-\sqrt{2}i\) compare this with a+bi and tell what's a and b ?

OpenStudy (zzr0ck3r):

((-1)-sqrt(2)i) *((-1)+sqrt(2)i) foil (-1)*(-1)+(1)*sqrt(2)i+(-1)sqrt(2)i-sqrt(2)^2*i*2 i^2 = -1

OpenStudy (anonymous):

what @hartnn said. forget about i, forget about minus signs, etc. it is just \(a^2+b^2\) a positive real number

OpenStudy (anonymous):

a= -1 and b=2?

hartnn (hartnn):

try b again.

hartnn (hartnn):

a is correct

OpenStudy (anonymous):

square root 2?

hartnn (hartnn):

\(b=-\sqrt{2}\) ok?

OpenStudy (anonymous):

oh ok. so u can have a negative? is that the product we found?

hartnn (hartnn):

the product is \(a^2+b^2\) can u calculate it?

OpenStudy (anonymous):

haha yes thanks!! you guys!

hartnn (hartnn):

welcome :) so what product u found?

OpenStudy (anonymous):

1+ -2 which = -1

hartnn (hartnn):

\((-\sqrt2)^2=2\) not -2.

OpenStudy (anonymous):

ait why? the negative doesnt get cancelled out

hartnn (hartnn):

because (-1)^2=1 so \((-\sqrt2)^2=(-1)^2(\sqrt{2}^2)=2\)

OpenStudy (anonymous):

ahhhh ok so the squared goes to both! thank you!!

hartnn (hartnn):

yup,so whats final answer?

OpenStudy (anonymous):

3

hartnn (hartnn):

correct :)

OpenStudy (anonymous):

thanks!!